Ice from a 3' by 4' by 5' box is being placed in paper cones. The cones are 3" in diameter and 5" tall. How many cones will be filled?
Answer by Cromlix(4381)

Hi there,
Volume of box = length x breadth x height
Volume of the box in inches:-
36 x 48 x 60 = 103680 inches^3
Volume of cone = 1/3 x π x radius^2 x height
Volume of cone = 1/3 x π x 1.5^2 x 5
Volume of cone = 11.78 inches^3
103680/11.78 = 8801 cones are filled.
Hope this helps :-)